OpenStudy (anonymous):

find the slope of a line through the points (2,5)and (3,5)

jimthompson5910 (jim_thompson5910):

Find the difference in the y coordinates. In other words, subtract the y coordinates What do you get when you do this?

OpenStudy (lalaly):

\[\large{\underline{slope}=\frac{y_2-y_1}{x_2-x_1}}\]

OpenStudy (anonymous):

jim i don't get it could show me the steps pls?

jimthompson5910 (jim_thompson5910):

The difference between the y coordinates is 5-5 = 0 The difference between the x coordinates is 2-3 = -1 Note: I'm subtracting the second coordinate from the first Now divide the y difference by the x difference: 0/(-1) = 0 So the slope is 0

OpenStudy (anonymous):

thnx but how do i do it in dy/dx form using first derivative

jimthompson5910 (jim_thompson5910):

you must be thinking of a totally different problem since you don't need to use derivatives in this simple problem

jimthompson5910 (jim_thompson5910):

derivatives do involve slopes, but to find the slope, you don't need to find the derivative

OpenStudy (anonymous):

ok how do i put it in dy/dx form

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean, but please don't confuse the two problems.

OpenStudy (anonymous):

am confused already

OpenStudy (anonymous):

y=x^4+5x^3-4x^2+7x-2

jimthompson5910 (jim_thompson5910):

yes this is a separate problem entirely, please post it as it's own thing.

OpenStudy (anonymous):

k don't you have any answer to it

jimthompson5910 (jim_thompson5910):

I do have an answer, just please post this as its own question